Issue with relatively similar CPS from 7800

I have inherited a 7800 from a new job, when I run the standard method for the technique all of the counts come out relatively similar. Running a blank and then a calibration standard returns the same results, approximately. The instrument has been cleaned from the lens all the way back to replacing the sampling probe in the last couple of days. However, I still get relatively the same number for CPS of every element. I attempted to check each of our solutions via the signal monitor, only paying attention to Li. When I run the blank it returns a count of ~150 for Li and when I run the standard I get ~250-300 for Li, using the signal monitor. Along with cleaning everything, I also switched out all the standards for ones ran on a 7850 that came back fine on that instrument. I replaced the ISIS 3 the other day because it wasn't drawing up correctly and I just noticed that an error: 2607 Warning Y-axis movement disabled due to a partially closed top cover today. I am trying to figure out why this instrument returns the same values for all elements. They have not run this instrument for a couple of months and last batch I see, seems to have worked fine. Any suggestions or ideas would be helpful, I appreciate it.

  • Those voltages look reasonable for the EM setting.  He gas is not used in startup.  Do they meet the values Chelsea listed above on the performance report?  If there is air in the He line, then counts will be very low until the He line is fully purged.  Just turning off the He gas in that mode will not give great sensitivity either, some of the other settings are much different than true no-gas mode.  If the performance report is giving good sensitivity and it is just your batch tune that is low counts, then the purge should help.

  • After I purged the He and ran a performance report which all passed. However, looking at my counts with signal monitor running gives me Li ~5k, Y ~0, and Tl ~900. Should I let the He purge overnight?

  • The performance report tells you that at least your sample introduction is in good condition since as Tom mentioned it does not utilize Helium mode. This is good news since it confirms you don’t have an issue with your nebulizer, cones, lenses, etc

    In Mass Hunter you can only purge a max of 12mL/min of He gas which is very little especially if your tank is farther away and it can take many hours. You can do it overnight if you like. What I will do instead to speed up the process is (with plasma off) disconnect the helium connection at the back of the ICP-MS and let the helium bleed for 5-10 minutes.

    Once I’ve done that, with the tank still open and has still bleeding, reconnect it to the ICP-MS. You’ve now purged everything up until the back of the instrument. Then go into Mass Hunter and set the flow to 12 and let it purge out another 5-10min. I often do this in signal monitor with the plasma on, the 205 counts should start to increase a little. This purges between the back of the instrument to the cell. You can return the flow back to the method parameters (often ~4-4.5mL) to confirm signal is improved.

    For testing purposes I would create a new batch from a preset method and choose one with low matrix settings. Once you get the helium purged out, perform an autotune with 1ppb tune solution and see if it looks okay. Then open your method batch and start comparing the settings, you can even manually input the lens settings if yours are very different and monitoring signal the whole time.

    If you are utilizing HMI (uses dilution gas) then 1ppb will not be adequate for tuning since HMI is less sensitive and I recommend changing to either a 10ppb or for very high HMI settings, 100ppb tune solution. 

  • I purged the He line and let it flow at 8 mL/min for around 4 hours then I then  turned on plasma and He then to about 5 mL/min. The 7800 does have HMI so I used the 10 ug/L tuning solution but am still seeing low counts overall. Autotune couldn't complete because counts were so low. It still does fine with a performance report but it seems that 205 is measuring a peak to left of the actual peak in the report and when I use the signal monitor to check counts 89 does have any counts now.

    I'm going to let it purge overnight at 2 mL/min to see if maybe it was not purged enough this morning, any other troubleshooting information to check out would be greatly appreciated.

  • At this point I don't think the helium is the issue if the 205 peak axis has shifted.

    What is the peak axis reading? It should be +/- 0.1amu of the assigned mass. Are you able to send a photo of both your performance report and batch tune report? Please make sure the photos show the entire report contents.

    If you have Pb contamination in the system, you will see a very large 206 peak next to 205. Any screenshots you can share will help tremendously!

  • This is the performance report from earlier today, after running He through it last night. I have not been able to get an autotune to complete since last week, it errors out saying counts are too low. The peak axis are all saying within 0.05 of expected.

Was this helpful?